Gdeinward:

Guys I really need help finding factors of 66 that add up to 15. I need two

Ferredoxin4:

Theres no such combination

Ferredoxin4:

Are you doing simplifying quadratics?

Gdeinward:

sort of

Gdeinward:

Im trying to solve \[\sqrt{x-7 } + 8 = x\]

Gdeinward:

not 7 but 2

Gdeinward:

so far ive gotten t0 -2 = x^2+ 15x + 66

Ferredoxin4:

Ah First subtract 8 Then square to remove the sqrt Youll get Then add 2 and subtract x to give a quadratic on 1 side and 0 on the other. Last you would solve. Just use the quadratic formula Youll get 2 values of x, plug in each to ensure it isnt extraneous

Gdeinward:

you must square both sides of the eqation.

Gdeinward:

I did all that. heres my steps below: \[\sqrt{x-2} = x+8\] x-2 = (x + 8) (x+8) x-2 (x^2 + 8x + 8x + 64

Gdeinward:

-2 = x^2 + 15 + 64

Gdeinward:

0 = x^2 + 15x + 66

Ferredoxin4:

Yeah so use the quadratic formula..do you know how to do it?

Gdeinward:

I know the formula ax^2 + bx + c = 0

Gdeinward:

but I need to factor the formula and to do that I must find two factors that ad up to the middle term

Gdeinward:

of the third term